OpenStudy (anonymous):

Given Quadrilateral ABCD and the points A(-2, 5), B(3, 5), C(3, -3) and D(-2, -3). Use the distance formula to find the lengths of each side. Give the most specific name for this polygon.

OpenStudy (anonymous):

ab=redical((3+2)^2+(5-5)^2)=5 bc=redical((3-3)^2+(-3-5)^2)=8 cd=redical((-2-3)^2+(-3+3)^2)=5 da=redical((-2+2)^2+(-3-5)^2)=8

OpenStudy (anonymous):

rectangle